# Attention on flow control: transformer-based reinforcement learning for lift regulation in highly disturbed flows

### v2

@@ I Introduction
An interest in large manned air vehicles has been extended to encompass unmanned air vehicles (UAVs) in recent years, particularly to serve roles in intelligence, surveillance, and delivery [1]. When performing a wide range of operations, it is expected for UAVs to achieve stable flight even in extreme environments where they may encounter sizable atmospheric disturbances (gusts) [2]. However, a large force and moment transient experienced during gust encounters may impact the flight path and safety of the aircraft and, over time, cause structural fatigue to the vehicle [3]. Therefore, many recent studies have been devoted to investigating the underlying unsteady aerodynamic physics of gust encounters [4, 5, 6] and designing control strategies to mitigate gust effect [7, 8, 9].
@@
Each combination of the control objectives (e.g., separation prevention, lift enhancement, or drag reduction), actuation methods (e.g., pitching, control surface, or synthetic jets), and the availability of measurements (e.g., pressure, velocity, acceleration, or force and moment), presents unique challenges and requires a tailored design approach. The control problem is manageable when gusts are of small amplitude, so that one can rely on linear control theory designed on a linear model of the aerodynamics. For example, Sedky et al. [8] designed a controller to mitigate the lift variation on an airfoil by modeling the lift response to pitching kinematics and transverse gust from the superposition of the classical Küssner model [10] and Wagner model [11]. Brunton and Rowley [12] developed a empirical state-space form of Theodorsen model [13] and leveraged it to design a robust controller to track a reference lift profile with $H_{\infty}$ loop-shaping approach. However, the use of classical aerodynamic models relies strictly on the fully attached flow assumption, and linear descriptions of the aerodynamics tend to fail if flow separation is induced, as is prone to occur when the disturbance is of large amplitude. In such circumstances one can no longer rely on superposition of indicial disturbance responses from sequences of gusts. Linear control tends to show degraded performance or even fail in these strong-gust conditions. In the absence of linear superposition, the control task is further confounded by the vast parameter space of the gusts and the strong nonlinear interactions with the wings [14, 5].
@@
Therefore, this challenge has spurred a large number of studies on modeling the nonlinear flow response and designing the appropriate control strategies when the flow is not fully attached. One popular approach is to linearize the flow around an equilibrium condition [15, 16, 17]. However, the limitation of this approach is that the model would be accurate only when the flow condition is near the equilibrium condition. Therefore, a gain-scheduling approach can be used to construct a series of linear models based on a series of equilibrium points. For example, Dawson et al. [18] constructed $8$ linear models for each corresponding to an incremental range of angle of attack, and these models could be switched during random pitch motion. However, they found that switching between the models leads to some degradation of accuracy when the pitching rate becomes relatively large. To address this problem, some other approaches have been pursued that incorporate proxy parameters of the relevant fluid flow physics into the system equations, such as linear parameter-varying model [19], Goman-Khrabrov model [20, 21], and the time delay and decay model [22]. Although these models show improved performance, their effectiveness depends heavily on the appropriate selection and fine-tuning of proxy parameters, which raises concerns about their generalizability in the vast parameter space of gust disturbances [14] and the diversity of available actuation mechanisms. Therefore, the development of a more generalizable and adaptive control strategy is still an open and compelling challenge.
@@
Fortunately, the development of machine learning and data science potentially provides us with data-driven tools to better model and control these complex flow problems. Among them, reinforcement learning (RL)—particularly when outfitted with a deep network for nonlinear function approximation, commonly referred to as deep reinforcement learning—has demonstrated notable success in a variety of flow control tasks. The examples include exploring effective swimming strategies [23, 24], learning an optimized foil flapping motion [25], reducing drag of a bluff body [26, 27, 28, 29], mitigating lift variation of an airfoil [30, 31, 32], and many others.
@@
Previous attempts to mitigate highly disturbed flows with RL made important progress, but each has left key gaps. In a flow context similar to the present paper, Beckers and Eldredge [31] trained a model-free agent in viscous, two-dimensional gust encounters for lift regulation of an airfoil using sparse surface pressures. Their policy and value networks were multi-layer perceptrons that processed a fixed history of observations. Such feed-forward models do not intrinsically model temporal dependencies, so the policy learned in a sequence of two gusts struggled when it was applied to long sequences of gusts. Furthermore, the learning curve converged at $5000$ episodes, which is prohibitively expensive from the perspective of CFD simulations. Therefore, for the purpose of reducing training cost, Liu et al. [30] proposed a model-based approach that learns a reduced-order surrogate model and optimizes the policy within the reduced space, which saved training cost significantly but must be refit when the disturbance class changes (e.g., the number of gusts), hence limiting generalization.
@@
We address both limitations in this work. We make use of the machine learning concept of self-attention and adopt the transformer [33], the most widely-used machine learning architecture that exploits this concept to model temporal dependencies, identifies correlations between different (and possibly widely separated) parts of the input observation sequence to approximate an output. We use the transformer output to approximate a belief state, which is used in turn to choose a corresponding control action via the proximal policy optimization (PPO) algorithm [34]. By this approach, we demonstrate that the policy learned in a small sequence of gusts generalizes well in long sequences. Although the transformer has been successfully applied to some fluid dynamics problems in recent years [25, 35], we note that, prior to this study, it was still unknown if it could be applied to the highly-disturbed flows, due to the unique challenges we discussed above. Furthermore, to reduce the computational cost, we leverage the techniques of pretraining and task-level transfer learning. The pretraining relies on the expert data from a well-designed proportional control (P control) strategy. We find that the RL control strategy that emerges outperforms the P control baseline, with an increasingly pronounced advantage as the number of gusts grows. The task-level transfer learning takes the RL agent trained in a single-gust environment as the starting point for the training in a multi-gust environment. We demonstrate that both these so-called warm-start techniques significantly speed up convergence compared to training from scratch.
@@
Additionally, both the previous works [30, 31] only explored a mid-chord pitching configuration. To investigate how the location of the pivot point affects the control performance of the RL agent, we revisit the pivot configuration and find that a quarter-chord pitching strategy spends much less control effort while achieving the lift regulation better than in mid-chord pitching. By analyzing the vorticity field, pressure distribution, lift history, and the corresponding pitching history, we offer physical insights into the superiority of quarter-chord pitching control.
@@ II.1 Flow control problem description
The flow environment is simulated computationally with a viscous incompressible flow solver based on the immersed boundary projection method [36]. The flow solver employs a vorticity-streamfunction formulation and utilizes a lattice Green’s function to solve the streamfunction Poisson equation, allowing for a more compact domain compared to traditional computational aerodynamics studies. From here on, we only report the time in convective time units, $t^{*}=tU_{\infty}/c$, and the spatial coordinates in chord lengths, $x^{*}=x/c$ and $y^{*}=y/c$. The computational domain spans [1.395,2.175][-1.395,2.175] in the $x^{*}$ direction and [0.885,0.885][-0.885,0.885] in the $y^{*}$ direction and employs free-space conditions on all boundaries. In this computational domain setup, the vortex gust and the resulting airfoil response are ensured to remain between the upper and lower boundaries and freely convect through the downstream boundary. We employ a uniform Cartesian grid with grid size $\Delta x^{*}=0.015$ and a uniform time step size $\Delta t^{*}=0.006$. Each simulation case lasts for $3$ convective time units, but is truncated early if $|\alpha|>\pi/4$ to avoid unrealistic configurations.
@@ II.2 Reinforcement learning framework
the basic RL framework, consisting of the environment and agent, is expressed as a Markov decision process (MDP). A MDP models the decision-making process by a tuple $\{\bm{S},\bm{A},\mathcal{T},\mathcal{R},\gamma\}$ where $\bm{S}$ is the state space, $\bm{A}$ is the action space, $\mathcal{T}$ is the transition probability function, $\mathcal{R}$ is the reward function, and $\gamma$ is the discounted factor: the agent receives a full state $s\in\bm{S}$ and selects an action $a\in\bm{A}$, the state $s$ is transitioned to a new state $s^{\prime}$ according to the transition probability function $\mathcal{T}(s^{\prime}|s,a)$, and a reward is also generated from $r=\mathcal{R}(s,a)$. However, the full state $s$ is not accessible in many real world systems, particularly in fluid dynamics, where this state includes the entire flow field. Instead, we only have partial information about the full state $s$ from available sensors, which leads to its formulation as a partial observable Markov decision process (POMDP) [37, 38]. A POMDP extends the MDP tuple to $\{\bm{S},\bm{A},\bm{O},\mathcal{T},\mathcal{Z},\mathcal{R},\gamma\}$ where, in addition to the MDP components, $\bm{O}$ is the observation space and $\mathcal{Z}=\mathcal{Z}(o^{\prime}|s,a)$ is the observation function, representing the probability of receiving observation $o^{\prime}$ given the state $s$ and action $a$. Therefore, instead of making a decision based on full state $s$ as in a MDP, the agent in a POMDP makes a decision based on a belief state $b(s)$, which represents a probability distribution over all possible states given the agent’s observation history. The agent learns a policy $\pi=\pi(a|b(s))$—that is, a policy conditioned on the distribution of possible states—that maximizes the discounted return
@@
As we discussed above, the agent relies on a belief state $b(s)$ to make a decision in a POMDP, which requires past observed information. Some RL algorithms propagate this belief state forward in time by folding in observations, and then condition their actions based on the most probable state. This approach is challenging in a fluid dynamics environment, where the state is high dimensional, though there has been recent success in sequentially estimating lower-dimensional representations of the flow state [39]. Alternatively, in fluid dynamics, some studies have directly employed the past observation history as a reduced representation of the state to approximate a POMDP as a MDP [31, 28, 35]. Francois-Lavet et al. [40] have shown that the agent can learn to encode the relevant past observations into a latent representation that serves as a surrogate of the probable state and make a decision based on this latent representation. Therefore, sequence models, such as recurrent neural networks (RNNs) [41] and transformers [33], play an important role since they can extract meaningful information from a time-dependent history sequence. One form of RNN, the long-short term memory (LSTM) [42], has been successfully applied to many complex fluid systems to model time-dependent behavior [30, 43, 44]. The mechanism of the LSTM is to update a hidden state at each time step based on the previous hidden state information and the new observation. This approach has two limitations: it is difficult to retain the information from distinct past observations when the sequence is very long, and the recursive process makes it impossible to process the sequence in parallel.
The limitations of the RNN motivate the recent popularity of the transformer [33], which leverages the self-attention mechanism to model a time sequence simultaneously. Specifically, the transformer first embeds the observation sequence into a higher dimensional space of tokens using a linear projection and a positional encoding (to ensure that temporal information is included):
@@
and represents the contextualized representation of the $i$th token. This output of the self-attention layer is fed into the feedforward network layer, which introduces nonlinearity to refine the token representation. The combination of the self-attention layer and the feedforward layer is an encoder layer. In the current work, our transformer employs two encoder layers with two attention heads in each layer. After the observation sequence $\left(o_{t-(N-1)},o_{t-(N-2)},\dots,o_{t}\right)$ is processed by the two-layer transformer structure, we obtain a sequence of contextualized representations $\left(\tilde{h}_{t-(N-1)},\tilde{h}_{t-(N-2)},\dots,\tilde{h}_{t}\right)$, where $\tilde{h}_{t}\in\mathbb{R}^{d_{\text{model}}}$ incorporates information from the entire observation window. In the current work, we extract the final element $\tilde{h}_{t}$ from this sequence and treat it as the latent representation of the probable state. This latent representation is then passed to the policy and value networks, which provide the probability distribution over the action space (for action selection) and the approximated expected return on that state (for state-value estimation), respectively. Fig. 3(b) summarizes this framework. For clarity, $d_{\text{obs}}=12$, $d_{\text{model}}=32$, $d_{k}=16$, and $n_{\text{head}}=2$ in this study.
We use the proximal policy optimization (PPO) algorithm [34] to approximate the optimal policy. Both the policy and the value networks consist of two hidden layers with $32$ hidden units each, where each hidden layer employs a Rectified Linear Unit activation function. The loss function that training seeks to minimize comprises three components. PPO, as a gradient-based model-free algorithm following the actor-critic framework, introduces a clipped surrogate objective to ensure updating within a safe region, constituting the first loss component
@@
where $\rho_{t}=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$ is the ratio of updated and old policies; $A_{t}$ is the advantage function, which helps the agent evaluate how good the action $a_{t}$ is at state $s_{t}$ compared to the expected value of this state $V(s_{t})$; and $\epsilon$ is the clip hyperparameter that controls the size of the safe region. The advantage function is estimated from the generalized advantage estimation (GAE) [45], defined as
@@
From previous work [39], it has been shown that surface pressure can be informed effectively by the change of flow field, and thus we incorporate surface pressure into the observation function. An observation window with a size of $N=20$ is employed throughout this study. Since we aim to find an optimal policy that enables the agent to track a reference lift profile, we design the reward function as follows:
@@
Before we complete this description of the methodology, we described one more technique that we use in this paper to accelerate the training. Finding a good expert policy from which to initiate training is often nontrivial. Therefore, we leverage task-level transfer learning [46] as another strategy when the expert data for the pretraining is not available. Transfer learning has been successfully applied to fluid dynamics problems recently [47, 29]. As depicted in Fig. 5, the RL agent is first trained in a single-gust environment (the source task), and once it acquires sufficient knowledge, the learned policy and value networks are transferred to a multi-gust environment (the target task), thereby accelerating learning and reducing training cost. This approach facilitates the application of RL to environments in which training is even more expensive: by first training a RL agent in a relatively easy or simplified environment, we obtain a well-informed initialization that serves as a warm start in the subsequent RL training in a more complex environment, and thus the overall training cost may be reduced.

### v3

@@ I Introduction
An interest in large manned air vehicles has been extended to encompass unmanned air vehicles (UAVs) in recent years, particularly to serve roles in intelligence, surveillance, and delivery [LAGHARI20238]. When performing a wide range of operations, it is expected for UAVs to achieve stable flight even in extreme environments where they may encounter sizable atmospheric disturbances (gusts) [Fukami2023]. However, a large force and moment transient experienced during gust encounters may impact the flight path and safety of the aircraft and, over time, cause structural fatigue to the vehicle [Wu2019]. Therefore, many recent studies have been devoted to investigating the underlying unsteady aerodynamic physics of gust encounters [Eldredge_Jones_2019, Jones2022, Narayanan2024] and designing control strategies to mitigate gust effect [Oduyela2014, Sedky_2022, Kerstens2011].
@@
Each combination of the control objectives (e.g., separation prevention, lift enhancement, or drag reduction), actuation methods (e.g., pitching, control surface, or synthetic jets), and the availability of measurements (e.g., pressure, velocity, acceleration, or force and moment), presents unique challenges and requires a tailored design approach. The control problem is manageable when gusts are of small amplitude, so that one can rely on linear control theory designed on a linear model of the aerodynamics. For example, Sedky et al. [Sedky_2022] designed a controller to mitigate the lift variation on an airfoil by modeling the lift response to pitching kinematics and transverse gust from the superposition of the classical Küssner model [Kussner1932] and Wagner model [Wagner1925]. Brunton and Rowley [BRUNTON2013] developed a empirical state-space form of Theodorsen model [Theodorsen1949] and leveraged it to design a robust controller to track a reference lift profile with $H_{\infty}$ loop-shaping approach. However, the use of classical aerodynamic models relies strictly on the fully attached flow assumption, and linear descriptions of the aerodynamics tend to fail if flow separation is induced, as is prone to occur when the disturbance is of large amplitude. In such circumstances one can no longer rely on superposition of indicial disturbance responses from sequences of gusts. Linear control tends to show degraded performance or even fail in these strong-gust conditions. In the absence of linear superposition, the control task is further confounded by the vast parameter space of the gusts and the strong nonlinear interactions with the wings [Jones2020, Jones2022].
@@
Therefore, this challenge has spurred a large number of studies on modeling the nonlinear flow response and designing the appropriate control strategies when the flow is not fully attached. One popular approach is to linearize the flow around an equilibrium condition [Rowley2005, Ma2011, Brunton_Rowley_Williams_2013]. However, the limitation of this approach is that the model would be accurate only when the flow condition is near the equilibrium condition. Therefore, a gain-scheduling approach can be used to construct a series of linear models based on a series of equilibrium points. For example, Dawson et al. [Dawson2015] constructed $8$ linear models for each corresponding to an incremental range of angle of attack, and these models could be switched during random pitch motion. However, they found that switching between the models leads to some degradation of accuracy when the pitching rate becomes relatively large. To address this problem, some other approaches have been pursued that incorporate proxy parameters of the relevant fluid flow physics into the system equations, such as linear parameter-varying model [Hemati2017], Goman-Khrabrov model [Goman1994, Williams2018], and the time delay and decay model [An2016]. Although these models show improved performance, their effectiveness depends heavily on the appropriate selection and fine-tuning of proxy parameters, which raises concerns about their generalizability in the vast parameter space of gust disturbances [Jones2020] and the diversity of available actuation mechanisms. Therefore, the development of a more generalizable and adaptive control strategy is still an open and compelling challenge.
@@
Fortunately, the development of machine learning and data science potentially provides us with data-driven tools to better model and control these complex flow problems. Among them, reinforcement learning (RL)—particularly when outfitted with a deep network for nonlinear function approximation, commonly referred to as deep reinforcement learning—has demonstrated notable success in a variety of flow control tasks. The examples include exploring effective swimming strategies [Novati2017, Verma2018], learning an optimized foil flapping motion [Wang2024], reducing drag of a bluff body [Fan2020, Rabault2019, Xia2024, YAN2025], mitigating lift variation of an airfoil [Liu2025, Beckers2024, Renn2022], and many others.
@@
Previous attempts to mitigate highly disturbed flows with RL made important progress, but each has left key gaps. In a flow context similar to the present paper, Beckers and Eldredge [Beckers2024] trained a model-free agent in viscous, two-dimensional gust encounters for lift regulation of an airfoil using sparse surface pressures. Their policy and value networks were multi-layer perceptrons (MLP) that processed a fixed history of observations. Such feed-forward models do not intrinsically model temporal dependencies, so the policy learned in a sequence of two gusts struggled when it was applied to long sequences of gusts. Furthermore, the learning curve converged at $5000$ episodes, which is prohibitively expensive from the perspective of CFD simulations. Therefore, for the purpose of reducing training cost, Liu et al. [Liu2025] proposed a model-based approach that learns a reduced-order surrogate model and optimizes the policy within the reduced space, which saved training cost significantly but must be refit when the disturbance class changes (e.g., the number of gusts), hence limiting generalization.
@@
We address both limitations in this work. We make use of the machine learning concept of self-attention and adopt the transformer [Vaswani2017], the most widely-used machine learning architecture that exploits this concept to model temporal dependencies, identifies correlations between different (and possibly widely separated) parts of the input observation sequence to approximate an output. We use the transformer output to approximate a belief state, which is used in turn to choose a corresponding control action via the proximal policy optimization (PPO) algorithm [schulman2017]. By this approach, we demonstrate that the policy learned in a small sequence of gusts generalizes well in long sequences. Although the transformer has been successfully applied to some fluid dynamics problems in recent years [Wang2024, Weissenbacher2025], we note that, prior to this study, it was still unknown if it could be applied to the highly-disturbed flows, due to the unique challenges we discussed above. Furthermore, to reduce the computational cost, we leverage the techniques of pretraining and task-level transfer learning. The pretraining relies on the expert data from a well-designed proportional control (P control) strategy. We find that the RL control strategy that emerges outperforms the P control baseline, with an increasingly pronounced advantage as the number of gusts grows. The task-level transfer learning takes the RL agent trained in a single-gust environment as the starting point for the training in a multi-gust environment. We demonstrate that both these so-called warm-start techniques significantly speed up convergence compared to training from scratch.
@@
Additionally, both the previous works [Liu2025, Beckers2024] only explored a mid-chord pitching configuration. To investigate how the location of the pivot point affects the control performance of the RL agent, we revisit the pivot configuration and find that a quarter-chord pitching strategy spends much less control effort while achieving the lift regulation better than in mid-chord pitching. By analyzing the vorticity field, pressure distribution, lift history, and the corresponding pitching history, we offer physical insights into the superiority of quarter-chord pitching control.
@@ II.1 Flow control problem description
The flow environment is simulated computationally with a viscous incompressible flow solver based on the immersed boundary projection method [eldredge2022]. The flow solver employs a vorticity-streamfunction formulation and utilizes a lattice Green’s function to solve the streamfunction Poisson equation, allowing for a more compact domain compared to traditional computational aerodynamics studies. From here on, we only report the time in convective time units, $t^{*}=tU_{\infty}/c$, and the spatial coordinates in chord lengths, $x^{*}=x/c$ and $y^{*}=y/c$. The computational domain spans [1.395,2.175][-1.395,2.175] in the $x^{*}$ direction and [0.885,0.885][-0.885,0.885] in the $y^{*}$ direction and employs free-space conditions on all boundaries. In this computational domain setup, the vortex gust and the resulting airfoil response are ensured to remain between the upper and lower boundaries and freely convect through the downstream boundary. We employ a uniform Cartesian grid with grid size $\Delta x^{*}=0.015$ and a uniform time step size $\Delta t^{*}=0.006$. Each simulation case lasts for $3$ convective time units, but is truncated early if $|\alpha|>\pi/4$ to avoid unrealistic configurations.
@@ II.2 Reinforcement learning framework
the basic RL framework, consisting of the environment and agent, is expressed as a Markov decision process (MDP). A MDP models the decision-making process by a tuple $\{\bm{S},\bm{A},\mathcal{T},\mathcal{R},\gamma\}$ where $\bm{S}$ is the state space, $\bm{A}$ is the action space, $\mathcal{T}$ is the transition probability function, $\mathcal{R}$ is the reward function, and $\gamma$ is the discounted factor: the agent receives a full state $s\in\bm{S}$ and selects an action $a\in\bm{A}$, the state $s$ is transitioned to a new state $s^{\prime}$ according to the transition probability function $\mathcal{T}(s^{\prime}|s,a)$, and a reward is also generated from $r=\mathcal{R}(s,a)$. However, the full state $s$ is not accessible in many real world systems, particularly in fluid dynamics, where this state includes the entire flow field. Instead, we only have partial information about the full state $s$ from available sensors, which leads to its formulation as a partial observable Markov decision process (POMDP) [ASTROM1965, KAELBLING1998]. A POMDP extends the MDP tuple to $\{\bm{S},\bm{A},\bm{O},\mathcal{T},\mathcal{Z},\mathcal{R},\gamma\}$ where, in addition to the MDP components, $\bm{O}$ is the observation space and $\mathcal{Z}=\mathcal{Z}(o^{\prime}|s,a)$ is the observation function, representing the probability of receiving observation $o^{\prime}$ given the state $s$ and action $a$. Therefore, instead of making a decision based on full state $s$ as in a MDP, the agent in a POMDP makes a decision based on a belief state $b(s)$, which represents a probability distribution over all possible states given the agent’s observation history. The agent learns a policy $\pi=\pi(a|b(s))$—that is, a policy conditioned on the distribution of possible states—that maximizes the discounted return
@@
As we discussed above, the agent relies on a belief state $b(s)$ to make a decision in a POMDP, which requires past observed information. Some RL algorithms propagate this belief state forward in time by folding in observations, and then condition their actions based on the most probable state. This approach is challenging in a fluid dynamics environment, where the state is high dimensional, though there has been recent success in sequentially estimating lower-dimensional representations of the flow state [mousavi2025]. Alternatively, in fluid dynamics, some studies have directly employed the past observation history as a reduced representation of the state to approximate a POMDP as a MDP [Beckers2024, Xia2024, Weissenbacher2025]. Francois-Lavet et al. [francoislavet2019] have shown that the agent can learn to encode the relevant past observations into a latent representation that serves as a surrogate of the probable state and make a decision based on this latent representation. Therefore, sequence models, such as recurrent neural networks (RNNs) [rumelhart1986] and transformers [Vaswani2017], play an important role since they can extract meaningful information from a time-dependent history sequence. One form of RNN, the long-short term memory (LSTM) [Hochreiter1997], has been successfully applied to many complex fluid systems to model time-dependent behavior [Liu2025, Hou2019, Hasegawa2020]. The mechanism of the LSTM is to update a hidden state at each time step based on the previous hidden state information and the new observation. This approach has two limitations: it is difficult to retain the information from distinct past observations when the sequence is very long, and the recursive process makes it impossible to process the sequence in parallel.
The limitations of the RNN motivate the recent popularity of the transformer [Vaswani2017], which leverages the self-attention mechanism to model a time sequence simultaneously. Specifically, the transformer first embeds the observation sequence into a higher dimensional space of tokens using a linear projection and a positional encoding (to ensure that temporal information is included):
@@
and represents the contextualized representation of the $i$th token. This output of the self-attention layer is fed into the feedforward network layer, which introduces nonlinearity to refine the token representation. The combination of the self-attention layer and the feedforward layer is an encoder layer. In the current work, our transformer employs two encoder layers with two attention heads in each layer. After the observation sequence $\left(o_{t-(N-1)},o_{t-(N-2)},\dots,o_{t}\right)$ is processed by the two-layer transformer structure, we obtain a sequence of contextualized representations $\left(\tilde{h}_{t-(N-1)},\tilde{h}_{t-(N-2)},\dots,\tilde{h}_{t}\right)$, where $\tilde{h}_{t}\in\mathbb{R}^{d_{\text{model}}}$ incorporates information from the entire observation window. In the current work, we extract the final element $\tilde{h}_{t}$ from this sequence and treat it as the latent representation of the probable state. This latent representation is then passed to the policy and value networks, which provide the probability distribution over the action space (for action selection) and the approximated expected return on that state (for state-value estimation), respectively. Fig. 3(b) summarizes this framework. For clarity, $d_{\text{obs}}=12$, $d_{\text{model}}=32$, $d_{k}=16$, and $n_{\text{head}}=2$ in this study. We also note that the adopted transformer structure in the present study is empirically demonstrated as a practical choice but may not necessarily be the only choice, and a comprehensive architecture benchmark (e.g. transformer vs. LSTM vs. MLP) will be an interesting direction for future work.
We use the proximal policy optimization (PPO) algorithm [schulman2017] to approximate the optimal policy. Both the policy and the value networks consist of two hidden layers with $32$ hidden units each, where each hidden layer employs a Rectified Linear Unit activation function. The loss function that training seeks to minimize comprises three components. PPO, as a gradient-based model-free algorithm following the actor-critic framework, introduces a clipped surrogate objective to ensure updating within a safe region, constituting the first loss component
@@
where $\rho_{t}=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$ is the ratio of updated and old policies; $A_{t}$ is the advantage function, which helps the agent evaluate how good the action $a_{t}$ is at state $s_{t}$ compared to the expected value of this state $V(s_{t})$; and $\epsilon$ is the clip hyperparameter that controls the size of the safe region. The advantage function is estimated from the generalized advantage estimation (GAE) [schulman2018], defined as
@@
From previous work [mousavi2025], it has been shown that surface pressure can be informed effectively by the change of flow field, and thus we incorporate surface pressure into the observation function. An observation window with a size of $N=20$ is employed throughout this study. A comparison between the RL agents trained with observation window sizes $N=20$ and $N=40$, as shown in appendix A, suggests that $N=20$ is an efficient and sufficient choice in the present study. Since we aim to find an optimal policy that enables the agent to track a reference lift profile, we design the reward function as follows:
@@
Before we complete this description of the methodology, we described one more technique that we use in this paper to accelerate the training. Finding a good expert policy from which to initiate training is often nontrivial. Therefore, we leverage task-level transfer learning [bozinovski2020] as another strategy when the expert data for the pretraining is not available. Transfer learning has been successfully applied to fluid dynamics problems recently [Wang20222, YAN2025]. As depicted in Fig. 5, the RL agent is first trained in a single-gust environment (the source task), and once it acquires sufficient knowledge, the learned policy and value networks are transferred to a multi-gust environment (the target task), thereby accelerating learning and reducing training cost. This approach facilitates the application of RL to environments in which training is even more expensive: by first training a RL agent in a relatively easy or simplified environment, we obtain a well-informed initialization that serves as a warm start in the subsequent RL training in a more complex environment, and thus the overall training cost may be reduced.